OpenStudy (anonymous):

Volume of a soup can. d=5.3cm h=17.7cm

OpenStudy (anonymous):

= area of circle times the height area of circle = pi squared times half the diameter

OpenStudy (anonymous):

The volume of the soup inside will of course be a little less. ;-)

OpenStudy (anonymous):

I am still confused how to use that formula.

OpenStudy (anonymous):

Volume of can = area of circle * height of the can

OpenStudy (anonymous):

here can is a cylinder. Cylinder of like a pile of circles placed one upon other! Hope it makes some sense

OpenStudy (anonymous):

I used that formula but getting a different answer from what I have in the book.

OpenStudy (anonymous):

to calculate area of circle what formula are u using? pi*d square/4 ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

use that formula.. when radius is given it is pi*r square. since r = d/2 we use that

OpenStudy (anonymous):

if d=5.3cm, then radius = 2.65cm square the radius = 7.0225cm^2 multiply by pi = 22.061834384625cm^2 (used 3.14159265 for pi) multiply by height, h=17.7cm, --> 390.4944686... cm^3 which you can round to 390 or 390.5 since that is the limit of precision of your input values.

OpenStudy (anonymous):

General formula is A=pi*r^2 Make sure to use proper order of operations and don't round off any of your intermediate values. Only ever round off final answers.

OpenStudy (anonymous):

Oops, meant to also include general formula for cylinder volume, V=pi*r^2*h Remember to square the radius first and then multiply by pi and by the height.

OpenStudy (anonymous):

Wait there is no radius Cliffsedge its only diameter and height provided.

OpenStudy (anonymous):

Use the definitions from geometry: radius = half of diameter | diameter = twice the radius. If you don't want to use that substitution then use the formula provided by srinu7j: Acircle = pi*d^2/4 Vcylinder = Acircle*height

OpenStudy (anonymous):

Got it!

OpenStudy (anonymous):

Thank You! Cliffsedge and srinu7j